A time study analysis shows an assembly operation for 50 cycles with an average cycle time of 15.00 minutes. If you assigned a performance rating of .95 and a work-day allowance factor of 20 percent, answer the following question:

What is the Allowance Factor?
Answer #1

As we see, a work day allowance factor is 20 percent = 0.2

But there are two ways in which this will be used:

  • Standard Time = Normal Time / ( 1 - Allowance)
  • Observed Time = Standard Time / (Performance rating)(1 + Allowance)

Hence, there will be two allowance factors.

While calculating Standard Time, Allowance factor = 1 - 0.2 = 0.8

And for Observed Time, Allowance factor = 1 + 0.2 = 1.2